Jack is 3 times as old as Lacy. 4 years ago the sum of their ages was 12. How old are they now?
Answer by TimothyLamb(4379)

x = jack age now
y = lacy age now
---
x = 3y
(x - 4) + (y - 4) = 12
---
put the system of linear equations into standard form
---
x - 3y = 0
x + y - 8 = 12
---
x - 3y = 0
x + y = 20
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = jack age now = 15
y = lacy age now = 5
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php